b) Determine the wavelength for a transition from the  = 0 to the  = 1 level. Is this
transition in the IR region of the electromagnetic spectrum?
c) Which of the following molecules exhibit rotational and or vibrational spectra
H2, HF, CO, NO
1
Expert's answer
2018-04-27T14:19:08-0400
b) The photon energy in the transition from the level v = 0 to v = 1 is:
Eph = ∆E = E1 – E0 = 3/2hv0 – ½ hv0 = hv0;
Since the energy of a photon is given by Ephoton = hv, the frequency of the transition from the first to the next second vibrational level is:
Eph = hv;
λ = c/v0 = 3*108/8.963*1013 = 3.345 * 10-6 m = 3345 nm.
c) Hydrogen atom (H2) vibrational spectra;
HF show rotational or vibrational spectra, since others are linear molecules.
CO will show rotational spectra while H2 will no.